# Toward General-Purpose Robots via Foundation Models: A Survey and Meta-Analysis

### v2

@@ Abstract
Building general-purpose robots that can operate seamlessly, in any environment, with any object, and utilizing various skills to complete diverse tasks has been a long-standing goal in Artificial Intelligence. Unfortunately, however, most existing robotic systems have been constrained—having been designed for specific tasks, trained on specific datasets, and deployed within specific environments. These systems usually require extensively-labeled data, rely on task-specific models, have numerous generalization issues when deployed in real-world scenarios, and struggle to remain robust to distribution shifts. Motivated by the impressive open-set performance and content generation capabilities of web-scale, large-capacity pre-trained models (i.e., foundation models) in research fields such as Natural Language Processing (NLP) and Computer Vision (CV), we devote this survey to exploring (i) how these existing foundation models from NLP and CV can be applied to the field of robotics, and also exploring (ii) what a robotics-specific foundation model would look like. We begin by providing an overview of what constitutes a conventional robotic system and the fundamental barriers to making it universally applicable. Next, we establish a taxonomy to discuss current work exploring ways to leverage existing foundation models for robotics and develop ones catered to robotics. Finally, we discuss key challenges and promising future directions in using foundation models for enabling general-purpose robotic systems. We encourage readers to view our living GitHub repository [^5] The current version of this paper is v1.1-2023.12 (In the format of ‘[major].[minor]-YYYY.MM’). of resources, including papers reviewed in this survey as well as related projects and repositories for developing foundation models for robotics: https://robotics-fm-survey.github.io/.
@@ 1.1 Introduction
We still face many challenges in developing autonomous robotic systems that can operate in and adapt to different environments. Previous robotic perception systems that leverage conventional deep learning methods usually require a large amount of labelled data to train the supervised learning models [1, 2, 3]; meanwhile, the crowdsourced labelling processes for building these large datasets remains rather expensive. Moreover, due to the limited generalization ability of classical supervised learning approaches, the trained models usually require carefully-designed domain adaptation techniques to deploy these models to specific scenes or tasks [4, 5], which often require further steps of data-collection and labelling. Similarly, classical robot planning and control methods often require carefully modelling the world, the ego-agent’s dynamics, and/or other agents’ behavior [6, 7, 8]. These models are built for each individual environment or task and often need to be rebuilt as changes occur, exposing their limited transferability [8]; in fact, in many cases, building an effective model can be either too expensive or intractable. Although deep (reinforcement) learning-based motion planning [9, 10] and control methods [11, 12, 13, 14] could help mitigate these problems, they also still suffer from distribution shifts and reductions in generalizability [15, 16].
@@
Concurrent to the challenges faced in building generalizable robotic systems, we notice significant advances in the fields of Natural Language Processing (NLP) and Computer Vision (CV)—with the introduction of Large Language Models (LLMs) [17] for NLP, the introduction of high-fidelity image generation with diffusion models [18, 19], and zero-shot/few-shot generalization of CV tasks with large-capacity vision models and Vision Language Models (VLMs) [20, 21, 22]. Coined “foundation models” [23], or simply Large Pre-Trained Models (LPTMS), these large-capacity vision and language models have also been applied in the field of robotics [24, 25, 26], with the potential for endowing robotic systems with open-world perception, task planning, and even motion control capabilities. Beyond just applying existing vision and/or language foundation models in robotics, we also see considerable potential for the development of more robotics-specific models, e.g., the action model for manipulation [27, 28] or motion planning model for navigation [29]. These robotics foundation models show great generalization ability across different tasks and even embodiments. Vision/language foundation models have also been applied directly to robotic tasks [30, 31], showing the possibility of fusing different robotic modules into a single unified model.
@@
In this paper, we investigate where foundation models are leveraged within robotics, am aim to understand how foundation models could help mitigate core robotics challenges. We use the term “foundation models for robotics” to include two distinct aspects: (1) the application of existing (mainly) vision and language models to robotics, largely through zero-shot and in-context learning; and (2) developing and leveraging robotics foundation models specially for robotic tasks by using robot-generated data. We summarize the methodologies of foundation models for robotics papers and conduct a meta-analysis of the experimental results of the papers we surveyed. A summary of the major components of this paper in Figure 4.
Figure 2: Overall structure of this survey paper. The left side shows the time dimension governing the emergence of foundation models, and the right side shows the logical induction, with respect to applying and developing foundation models for robotics. Sections 2 and 3 answer the “what” questions: what is robotics, what are the challenges in robotic problems, and what are foundation models. Section 4 and 5 deal with the “why” and “how” questions: why do we need foundation models in robotics, how can foundation models be applied in robotics, and how do foundation models specially designed for robotics work. Section 6 closes with existing work, and posits what future models might look like.
@@ 1.2 Related Survey Papers
Recently, with the popularity of foundation models, there are various survey papers on vision and language foundation models that are worth mentioning [32, 33, 34, 35]. These survey papers cover foundation models, including Vision Foundation Models (VFMs) [36, 37], Large Language Models (LLMs) [34], Vision-Language Models (VLMs) [38, 39], and Visual Content Generation Models (VGMs) [32]. Since foundation models for robotics are still relatively nascent areas, there are not so many existing survey papers combining foundation models and robotics, perhaps the most relevant survey papers are [35, 40, 41, 42, 43], however, there are still significant differences between those papers and ours, for instance: Yang et al. [35] and Wang et al. [40] focus on broadly-defined autonomous agents, instead of physical robots; Lin et al. [41] focus on LLMs for navigation; the connection between foundation models and robotics is limited in [42]. Compared with  [43], we propose more breakdown of current research methodologies, provide detailed analysis of the experiments, and also focus on how foundation models could resolve the typical robotic challenges. Concurrently, Firoozi et. al. [44] conducted a survey regarding foundation models in robotics. Both their and our works shed lights on the opportunities and challenges of using foundation models in robotics, and identifies key pitfalls to scale them further. Their work focuses on how foundation models contribute to improving robot capabilities, and challenges going forward. Comparatively, our survey attempts to taxonomize robotic capabilities together with foundation models to advance those capabilities. We also propose a dichotomy between robotic(-first) foundation models and other foundation models used in robotics, and provides a thorough meta-analysis of the papers we survey.
@@
2 Preliminaries
In this section, we walk through the preliminaries to help readers better understand the contents of this paper. Since we focus on foundation models centered around robotics, we will first introduce the basics of robotics and the current state of the art. These preliminaries will focus on the methods before foundation models were applied in robotics. For the ease of organization, we introduce robotic modules based on their functionalities, e.g., perception 2.1.1, planning 2.1.2 and control 2.1.3. We note that, although we introduce these modules separately, the boundaries between these modules are often blurred  [45, 12]: these modules are often synergistically connected, facilitating end-to-end differentiability—allowing gradients to flow across different modules, especially in learning-based approaches. The second focus of this section will be to provide an introduction to foundation models, mainly situated in the fields of NLP and CV; these models include: LLMs, VLMs, vision foundation models, as well as text-conditioned image generation models.
2.1 Ingredients of a Robotic System
2.1.1 Robot Perception
Robots require perceptual mechanisms, in order to extract semantic knowledge from raw sensor observations, establish state representations, and enact reasoning in their operating environments. Different from typical computer vision systems, robotic perception emphasizes real-time capability, the use of multiple modalities (RGB, depth, LiDAR, IMU, tactile, etc.), the coupling with other robotic system modules (decision making, planning, control), and grounding with the embodiment and the environment [46, 47].
Passive Perception
The most common use-case for typical computer vision algorithms in a robot perception system is for scene understanding. Here, the goal is to extract insights about the semantic and geometric properties of an environment by processing visual signals (e.g., 2D image data, RADAR information, LiDAR/RGB-D point clouds), perhaps to perform specific tasks like object detection & tracking, semantic segmentation, pose-estimation, novel view synthesis, or scene reconstruction [48, 49, 50, 51, 52]. However, the problem with the dominant learning-based approaches is that they primarily rely on large amounts of labeled data for training, where it is particularly challenging to obtain large-scale labels. Furthermore, these approaches tend to break down in out-of-distribution scenarios, rendering them too brittle for extensive deployment. As discussed in Section 2.2.1, this problem has been largely alleviated with the impressive open-set capabilities of visual foundation models [53, 22].
State Estimation
State estimation is the challenging problem of estimating the poses or velocities of robots based on sensor measurements. The task of Simultaneous Localization and Mapping (SLAM) seamlessly integrates the pose estimation and mapping problems together. State estimation and SLAM can be addressed using various sensor modalities, as evidenced by the collection of vision-based approaches [54, 55, 56, 57, 58, 59], LiDAR-based techniques [60, 61, 62], methods leveraging Inertial Measurement Units (IMU) [63, 64], and sensor fusion methods involving multiple sensors [65, 66, 67, 68, 69]. While traditional approaches typically rely on rigorous geometry-based solutions, more recently there has been a growing interest in learning-based approaches, which leverage supervised [70, 71] and self-supervised [72, 73, 74, 75] methods. These learning-based methods have demonstrated their ability to provide accurate pose-tracking results and, in some cases, achieve dense reconstruction even without the use of depth sensors.
Active Perception
The previously introduced approaches give a robot the ability to perceive the environment only in a passive manner, where information gain or decision-making does not play a role in how the perception system evolves, temporally. Since robots move and often interact with the environment, the robot should also be able to perceive the environment in an active manner. Prior approaches to the active perception problem have looked to the lens of interacting with the environment [76] or by changing the viewing direction to obtain more visual information [77, 78, 79].
2.1.2 Robot Decision-making and Planning
Classical Planning
Planning for robotics is a process of organizing a set of actions that a robot should execute, given a model of itself and the world, to reach desired states and minimize cumulative costs when executing those actions. Motion planning aims to find a collision-free path to reach desired states. Search-based planning [80, 81, 82, 83, 84] computes robots’ trajectories based on discrete representations of problems, taking advantage of heuristics and graphs. Another major area, sampling-based planning [85, 86, 87, 88, 89, 90], seeks to randomly sample points in configuration spaces, find paths to desired states by connecting nearby points or incrementally generating next states, and is well-suited for high-dimensional planning and continuous spaces.
Task planning [91] deals with discrete domains with tractable and compact representations for large state-spaces, usually exploiting structural properties of the domains and object-level abstractions. The approaches for compact representations include factoring the state representations into a set of smaller state variables, use of a set of preconditions that specify the states in which robots can execute actions, and object-oriented abstraction for symbolic reasoning [92].
Learning-based planning
There exist newer works in planning that use reinforcement learning [93, 94] and formulate motion-planning as an end-to-end problem [9, 10]. In navigation, ReViND [95] and FastRLAP [94] use offline reinforcement learning to learn the planning policy for visual navigation. By optimizing a value function from a static dataset, robots could learn driving behaviors in a short period of time [94]. In addition, the advantage of utilizing reward re-labeling allows the robot to earn different navigation behavior according to the reward specified by the user [95]. Learning for task planning includes recent works on integrating reinforcement learning with task planning [96, 97] to improve adaptability in dynamic environments and generate plans, and works on learning symbolic abstract model and representations for task planning [98, 99].
2.1.3 Robot Action Generation
Classical Control
Low-level action control, achieved via direct actuation or motor control, is the last step in most robotics stacks. This part of the stack is usually dependent on the exact platform and often incorporates dynamics and actuator constraints, thereby ensuring the feasibility of the generated action by keeping the robot within its operational envelope. While the control input is usually in continuous space, motion primitives are often used to provide a discrete set of actions for ease of interfacing with higher-level decision loops. Arguably, PID control loops are the most widely used lower-level control structures for robotic systems. When a cost function is available, optimization-based methods, a.k.a. optimal control, such as Model Predictive Control (MPC) and its variants [100, 101, 102, 103] are often used to generate action sequences in a receding-horizon setup. The Model Predictive Path Integral (MPPI) controller [104], a variant of MPC, is widely used in its sampling formulation on learned cost maps.
Learning-based Control
Applying imitation learning [105] or reinforcement learning [106] in robotic control has been studied for decades. With the success of deep learning [107] and deep reinforcement learning [108, 109], we see this line of research getting a good number of success stories in recent years [110, 12, 111, 112, 13, 113].
Imitation learning aims to learn a control policy by imitating demonstrations from some expert, which could be implicit in a dataset of trajectories. It can be in the form of supervised learning which directly learns actions from expert demonstrations [114], inverse reinforcement learning [115, 116] which learns reward functions, and adversarial imitation learning which learns the policy with generative adversarial networks [117, 118]. Imitation learning is widely used in various robot control applications, including: urban driving [101], high-speed car racing [119], autonomous drone acrobatics [110], learning locomotion skills by imitating animals [11], and quadrupedal agile skills via adversarial imitation learning [120].
Reinforcement learning (RL) [121] is typically leveraged in the context of a Markov Decision Process, in order to learn and optimize a control policy via accumulated rewards. Different from optimal control-based methods, RL methods may not require dynamics models. Many existing works in RL for robotics follow a model-free learning paradigm, wherein a policy learns to directly map sensor observations (e.g., images [12, 111, 122], proprioception [13, 113, 123], or both [124]) to generate actions. Model-free approaches usually have the drawback of being sample-inefficient. Model-based RL methods [125] provide a viable solution by learning a world dynamics model and then planning or learning the control policy. These world dynamics model can be in the form of visual observations [126, 127] or dynamics models based on proprioception [14]. However, aforementioned online RL methods in real-world may give rise to safely concerns [95]. Offline reinforcement learning approaches [95, 128, 129] attempt to remedy this problem, since the learning is based only on offline datasets. We have already seen a few works which apply offline RL in robotics, such as in visual navigation [93], high-speed ground vehicle driving [94], and manipulation [130, 131].
2.2 Introduction to Foundation Models
From the definition in [23], a foundation model is any model that is trained on broad data (generally using self-supervision at scale) that can be transferred or adapted (e.g., fine-tuned) to a wide range of downstream tasks. Existing successful foundation models are mainly from CV and NLP areas, e.g. (ranked from single modalities to multiple modalities): Vision Foundation Models (VFMs; Section 2.2.1), Large Language Models (LLMs; Section 2.2.3), Vision-Language Models (VLMs; Section 2.2.4), and Large Multimodal Models (LMMs; Section 2.2.5). Recently, we also see the rise of foundation models that are specifically designed for robotics tasks and trained on robotics data (“robotics foundation models”): this particular topic will be introduced, later, in Section 4.2.
2.2.1 Vision Foundation Models (VFMs)
Alongside the advent of LLMs and VLMs, several Vision-based Foundation Models (VFMs) have been proposed [132, 21, 20, 22]. Owing to their impressive domain-invariance and semantic properties at the pixel- and object-level [133, 134, 135, 136, 137], these vision foundation models have been widely adopted for downstream passive perception tasks. Furthermore, these major advances have been enabled either through self-supervision [133] and/or large-scale data curation [22, 21].
The family of self-supervised VFMs can be broadly organized into the following three subclasses: (1) Joint-Embedding Predictive Architectures (JEPA; [138]), (2) Contrastive Learning-based methods [53, 139], (3) and Masked Autoencoder (MAE; [132]) JEPAs employ a Bootstrap Your Own Latent (BYOL; [140]) style of self-supervision technique, where the primary supervisory signal is to predict similar embeddings across different augmentations of an image. Amongst the JEPA methods, the most notable ones are DINO [20], DINOv2 [22], I-JEPA [141] and MC-JEPA [142]. Recent explorations have shown that these joint-embedding-based approaches capture longer-range global patterns and shape-oriented features [133, 134]. On the other hand, contrastive learning-based methods leverage the weak supervision from multimodal data to learn a common latent space across different modalities. Notable methods include CLIP [53], which uses large-scale image-caption pairs. In addition to these two classes, MAEs [132] make up another class of models, trained to reconstruct masked inputs as a pretext task. Explorations have shown that these models capture local token-level semantic context, leading to their wide popularity for dense prediction problems such as semantic segmentation [133, 134].
Two notable VFMs that have been enabled by the careful curation of large datasets are the Segment Anything Model (SAM [21]) and DINOv2 [22]. SAM leveraged an iterative model prediction-based curation process to obtain 1 billion semantic segmentation masks for supervised learning. It has been showcased that the SAM models, trained on this large-scale curated data, show impressive instance segmentation performance across a wide range of domains. Similarly, DINOv2 [22] is a self-supervised model trained using model prediction-based curated data comprising 142 million images. It has been showcased that this large-scale self-supervision on curated data enables DINOv2 to perform better than task-specifically trained models and contrastive zero-shot models such as CLIP, while showcasing impressive semantic consistency.
2.2.2 Visual Content Generation Models (VGMs)
Text-conditioned image generation models have achieved great attention recently due to their astonishing ability to generate novel, high-fidelity images directly from language prompts, thanks to the progress in diffusion models [143]. GLIDE [144] is a text-conditioned diffusion model with both CLIP guidance and classifier-free guidance. DALLE-2 [18] proposes a two-stage diffusion model that consists of a prior that generates a CLIP image embedding given a text caption and a decoder that generates an image conditioned on the encoded image embedding. IMAGEN [19] is another text-conditioned diffusion model with classifier-free guidance. Different from previous approaches, it proposes dynamic thresholding to generate more photorealistic images and a U-Net structure to make the training more efficient. We name this type of foundation model as Visual Content Generation Models (VGMs) in this paper for convenience.
2.2.3 Large Language Models (LLMs)
A Large Language Model (LLM) is a type of language model notable for its ability to handle a variety of language tasks with minimal task-specific training data, setting it apart from conventional AI models [145]. The term large refers to both the model size and dataset size. Moreover, language signifies that the models are trained on an internet scale with a single modality, which is text. The key development in LLMs was the introduction of the transformer architecture, which allows for the efficient training of large-scale data due to the highly parallel nature of transformers, making the processing of extended text sequences more efficient. Two lines of work build upon the transformer architecture: the Generative Pre-trained Transformer (GPT) series [146, 17] and the Bidirectional Encoder Representations from Transformers (BERT) family [147]. GPT is trained as a decoder, with the task of predicting the next word in a sequence, whereas BERT is trained as an encoder, focusing on understanding the contextual relationships between sentences. However, according to Yang et al. [148], the BERT family has not seen significant advancements since 2021, with all current state-of-the-art LLMs building upon the GPT structure. Another technique that has contributed to the achievements of models like the GPT-3.5 model (ChatGPT) is the paradigm of Reinforcement Learning from Human Feedback (RLHF) [149]. This technique aligns the transformer’s output with human preferences learned through inverse reinforcement learning, enabling the model to produce more human-like and fluent sentences.
However, the limitations of LLMs are still significant, particularly in their tendency to hallucinate to produce plausible outputs. In addition, due to the limitations of being trained only on a single text modality, LLMs lack references to real-world objects and, therefore, cannot accurately answer questions such as, “Can I put this watermelon into the blender?”. This problem of lacking spatial grounding is especially important for the application of LLMs to Robotics. As a result, more recent research has started to integrate additional sources of information, such as vision, into transformer architectures.
2.2.4 Vision-Language Models (VLMs)
Vision-language models (VLMs) represent another significant advancement in generative models. These models excel in tasks that require understanding both visual content and language, such as open-set image classification [53], object detection [150], and segmentation [151], Visual Question Answering (VQA) [152], etc. These models leverage large-scale datasets and sophisticated neural network architectures, typically variants of transformers, to learn correlations between images and their textual descriptions or queries. This approach enables them to perform a range of tasks without task-specific training, showcasing impressive generalization abilities. They can be broadly categorized into two main groups based on their pre-training methods: contrastive learning models and generative pre-training models.
Contrastive learning based models, like CLIP [53] and ALIGN [153], are trained to understand the correspondence between images and texts by bringing the representations of matching image-text pairs closer in the embedding space while pushing non-matching pairs apart. CLIP, for instance, excels in a wide range of visual classification tasks with its ability to understand nuanced textual descriptions and their corresponding images. ALIGN, similarly, focuses on aligning large-scale image-text pairs, significantly improving performance on tasks like image captioning and visual question answering. Generative pre-training models, like ViLBERT [152] and VL-BEiT [154] take similar training approach as in LLMs 2.2.3.
Combing the strength of the both methods, Flamingo [155] incorporates a large frozen language model, retaining the in-context few-shot learning ability that is inherent to the pre-trained language model. Conversely, GIT [156] employs a large contrastively pre-trained image encoder, accompanied by a relatively small text decoder. Both Flamingo and GIT commence by pre-training an image encoder through contrastive learning, then subsequently undertaking generative pre-training.
More recently, BLIP [157] and BLIP-2 [158] have emerged as a transformative model in VLMs, introducing a curriculum learning strategy that bootstraps from simpler to more complex tasks, significantly enhancing performance in tasks like image captioning and visual question answering. The latest GPT iteration, GPT-4 [159], introduces the capability to process both textual and visual inputs. However, as of the last update, the technical details and the extent of these new multimodal capabilities are not released yet. Collectively, these models exemplify the rapid advancements in VLMs, each contributing to the robustness and adaptability of multimodal systems in understanding and in generating human-like responses based on visual data.
2.2.5 Large Multimodal Models (LMMs)
The combination of vision and language modalities reveal the huge potential of self-supervised learning. So it is natural to extend beyond vision and language to develop new types of foundation models with even more modalities. These models are coined Large Multimodal Models (LMMs). The additional modalities are rich and diverse, as in models that combine image, text, depth, thermal, and audio, e.g., ImageBind [139]; models combining text, image, video and audio, e.g., NExT-GPT [160] and Audio-GPT [161]; models combining language text and audio, e.g., SpeechGPT [162]; and models combining point cloud with vision and language, e.g., ULIP [163]. These LMMs use distinct training methods, such as contrastive pre-training  [139] or fine-tuning LLM by learning input/output projections [160]. The contrastive learning methods shown in these LMMs are fairly similar to LLMs, in the sense that embeddings of different modalities are aligned with each other. These multimodal contrastive learning-based methods enable cross-modal retrieval and bring in more interesting applications such as audio to image generation, etc.
3 Challenges in Robotics

### v3

@@ Abstract
Building general-purpose robots that operate seamlessly in any environment, with any object, and utilizing various skills to complete diverse tasks has been a long-standing goal in Artificial Intelligence. However, as a community, we have been constraining most robotic systems by designing them for specific tasks, training them on specific datasets, and deploying them within specific environments. These systems require extensively-labeled data and task-specific models. When deployed in real-world scenarios, such systems face several generalization issues and struggle to remain robust to distribution shifts. Motivated by the impressive open-set performance and content generation capabilities of web-scale, large-capacity pre-trained models (i.e., foundation models) in research fields such as Natural Language Processing (NLP) and Computer Vision (CV), we devote this survey to exploring (i) how these existing foundation models from NLP and CV can be applied to the field of general-purpose robotics, and also exploring (ii) what a robotics-specific foundation model would look like. We begin by providing a generalized formulation of how foundation models are used in robotics, and the fundamental barriers to making generalist robots universally applicable. Next, we establish a taxonomy to discuss current work exploring ways to leverage existing foundation models for robotics and develop ones catered to robotics. Finally, we discuss key challenges and promising future directions in using foundation models for enabling general-purpose robotic systems. We encourage readers to view our living GitHub repository  [^5] The current version of this paper is v2.1-2024.09 (In the format of ‘[major].[minor]-YYYY.MM’). of resources, including papers reviewed in this survey, as well as related projects and repositories for developing foundation models for robotics: https://robotics-fm-survey.github.io/.
@@ 1.1 Introduction
We face many challenges in developing general-purpose robotic systems that can operate in and adapt to different environments and perform multiple tasks. Previous robotic perception systems that leverage conventional deep learning methods usually require a large amount of labeled data to train the supervised learning models [1, 2, 3]; meanwhile, the crowdsourced labeling processes for building these large datasets remain rather expensive. Moreover, due to the limited generalization ability of classical supervised learning approaches, the trained models usually require carefully designed domain adaptation techniques to deploy these models to specific scenes or tasks [4, 5], which often require further steps of data-collection and labeling Similarly, classical robot planning and control methods often require carefully modeling the world, the ego-agent’s dynamics, and other agents’ behavior [6, 7, 8]. These models are built for each individual environment or task and often need to be rebuilt as changes occur, exposing their limited transferability [8]; in fact, building an effective model can be either too expensive or intractable. Although deep (reinforcement) learning-based motion planning [9, 10] and control methods [11, 12, 13, 14] could help mitigate these problems, they also still suffer from distribution shifts and reductions in generalizability [15, 16].
Figure 2: Overall structure of this survey paper. The left side shows the time dimension governing the emergence of foundation models, and the right side shows the logical induction, for applying and developing foundation models for robotics. Sections 2 and 3 answer the “what” questions: what are foundation models for robotics, what are the challenges in building general-purpose robots. Section 4 and 5 deal with the “why” and “how” questions: why do we need foundation models in general-purpose robotics, how can foundation models be applied in robotics, and how do foundation models specially designed for robotics work. Section 6 closes with existing work, and posits what future models might look like.
Concurrent to the challenges faced in building generalizable robotic systems, we notice significant advances in the fields of Natural Language Processing (NLP) and Computer Vision (CV)—with the introduction of Large Language Models (LLMs) [17] for NLP, diffusion models to generate high-fidelity images and videos [18, 19], and zero-shot/few-shot generalization of CV tasks with large-capacity Vision Foundation Models (VFM) [20, 21, 22], Vision Language Models (VLMs) and Large Multi-Modal Models (LMM) [23, 24] , coined foundation models [25].
@@
These large-capacity vision and language models have also been applied in the field of robotics [26, 27, 28], with the potential for endowing robotic systems with open-world perception, task planning, and even motion control capabilities. Beyond just applying existing vision and/or language foundation models in robotics, we also see considerable potential for the development of more robotics-specific models, e.g., the action model for manipulation [29, 30], motion planning model for navigation [31], and more generalist models which can conduct action generation though vision and language inputs hence Vision Language Action models (VLA) [32, 33, 30, 34]. These robotics foundation models show potential generalization ability across different tasks and even embodiments.
@@
In this paper, we investigate how foundation models serve as a potential solution for general-purpose robotics, aiming to understand how foundation models could help mitigate the core challenges general-purpose robotics face. We use the term “foundation models for robotics” to include two distinct aspects: (1) the application of existing (mainly) vision and language models to robotics, largely through zero-shot and in-context learning; and (2) developing and leveraging robotics foundation models especially for robotic tasks by using robot-generated data. We summarize the methodologies of foundation models for robotics papers and conduct a meta-analysis of the experimental results of the papers we surveyed. A summary of the major components of this paper in Figure 4.
@@ 1.2 Related Survey Papers
Recently, with the popularity of foundation models, there are various survey papers on vision and language foundation models that are worth mentioning [35, 36, 37, 38]. These survey papers cover foundation models, including Vision Foundation Models (VFMs) [39, 40], Large Language Models (LLMs) [37], Vision-Language Models (VLMs) [41, 42], and Visual Content Generation Models (VGMs) [35]. There are a few existing survey papers combining foundation models and robotics, perhaps the most relevant survey papers are [43, 38, 44, 45, 46, 47], however, there are still significant differences between those papers and ours, for instance: Yang et al. [38] and Wang et al. [44] focus on broadly-defined autonomous agents, instead of physical robots; Lin et al. [45] focus on LLMs for navigation; the connection between foundation models and robotics is limited in [46]. Compared with  [47], we propose a breakdown of current research methodologies, providing a detailed analysis of the experiments, and also focus on how foundation models could resolve the typical challenges for general-purpose robotics. Concurrently, Firoozi et al. [43] conducted a survey regarding foundation models in robotics. Both their and our works shed lights on the opportunities and challenges of using foundation models in robotics, and identifies key pitfalls to scale them further. Their work focuses on how foundation models contribute to improving robot capabilities, and challenges going forward. Comparatively, our survey attempts to taxonomize robotic capabilities together with foundation models to advance those capabilities. We also propose a dichotomy between robotic(-first) foundation models and other foundation models used in robotics, and provide a thorough meta-analysis of the papers we survey.
@@
2 Problem Formulation and Preliminaries
In this section, we summarize foundation models for robotics as a unified problem formulation. Let us define the foundation model relevant for robotics as a function $f$, which takes sensory inputs $\mathbf{x_{t}}$, context $\mathbf{c}$, and outputs $\mathbf{y_{t}}$ to fulfill the downstream robotic tasks.
$f(\mathbf{x_{tk},c_{k}})\rightarrow\mathbf{y_{tk}}\quad\quad\quad\forall k\in N\quad\forall t\in T$ (1)
Here $\mathbf{x}$ denotes sensory inputs like observations through visual cameras, textual scene descriptions, scene graphs, object poses and detections, audio signals through a microphone, haptic sensors, and so on. For multi-task policies, we need contextual information $\mathbf{c}$ that denotes task specification [30] or details on the embodiment [48].
The state $\mathbf{x}$ and context $\mathbf{c}$ are often defined in terms of visual observations and language instructions respectively. However, both $\mathbf{x}$ and $\mathbf{c}$ could be images, videos, language prompts, or a combination of multiple modalities. This also depends on how we define the scope of the environment.
The action $\mathbf{y}$ can be defined in terms of target objects’ pose in the map, task plans, next state, reward functions, and control outputs like target end-effector pose. If $\mathbf{y}$ is a task plan, we often assume that the low-level components of the plan can be executed perfectly. However, there are open questions on how to enable suitable and timely recovery in the robot policies when the low-level plan is not executed completely.
The data to train or prompt the model $f$ have several open questions. For training, how should expert demonstrations, offline data, or online experiences with apt reward functions be collected? For prompting, what instructions and examples enable desirable in-context learning? How should the models be trained and fine-tuned? In the next following sections, we aim to answer these questions by summarizing and analyzing existing research works.
3 Challenges on General-purpose Robots